OpenStudy (anonymous):

A drawer contains 6 red colored pencils and 2 pink colored pencils. What is the probability that you select two red colored pencils if you reach in the drawer and randomly grab two colored pencils without replacement? A) 15/28 B) 1/8 C) 1/12 D) 1/4 E) 1/6 Just completely blanked.

OpenStudy (klimenkov):

\(\frac{15}{28}\)

OpenStudy (anonymous):

How again. Normall'y this'd be easy for me but i've been doing too much other math. needa get back into SAT math things.

OpenStudy (anonymous):

\[\frac68\times\frac57=\frac{15}{56}\]

OpenStudy (anonymous):

K thx.